OpenStudy (anonymous):

How do I find the solution to a system? y=x^2+6x+6 y=4x+12

OpenStudy (radar):

Since y=y x^2 + 6x + 6 =4x + 12 simplify using algebra to: x^2 + 2x - 6 = 0 this should put you on familiar ground.........solve the quadratic then solve for y.

OpenStudy (radar):

I suggest completing the square method for solution of the quadratic. rearrange the quadratic for this method: x^2 + 2x = 6 x^2 + 2x + 1 = 6 + 1 (x + 1)^2 = 7 x + 1 = sq rt 7 or\[x + 1 = \sqrt{7}\] Please continue.